OpenStudy (anonymous):

If g(x) + xsin g(x) = x^2 , find g ' (0)

OpenStudy (anonymous):

\[g(x) + x \sin g(x) = x ^{2}\] find \[g ' (0) \]

OpenStudy (anonymous):

take the derivative and get \[g'(x)+\sin(g(x))+x\cos(g(x))g'(x)=2x\] now replace x by 0 and solve for \[g'(0)\]

OpenStudy (anonymous):

actually i am confused. is it \[g(x) + x \sin(x) g(x) = x ^{2}\] or \[g(x) + x \sin( g(x)) = x ^{2}\]?

OpenStudy (anonymous):

i am going to guess it is the first one, since i don't think you can do the second.

OpenStudy (anonymous):

you get \[g'(x)+\sin(x)g(x)+x\cos(x)g(x)+x\sin(x)g'(x)=2x\]now replace x by 0

OpenStudy (anonymous):

you get \[g'(0)=0\]

OpenStudy (anonymous):

.so its 2(0) so its 0?

OpenStudy (anonymous):

that is what i got, yes just 0

OpenStudy (anonymous):

thank u very much. btw can i ask u a question on teh chain rule?

OpenStudy (anonymous):

im confused when to use it and how?